What is the value of f(16)- f(0) when f(x)=4x–8 ?
Answer by stanbon(75887)

What is the value of f(16)- f(0) when f(x)=4x–8 ?
-------
f(16) = 4*16-8 = 64-8 = 56
f(0) = 4*0-8 = -8
------------------------------
f(16)-f(0) = 56 -(-8) = 64
----------------------
Chers,
Stan H.
---------------